# NoTeeline: Supporting Real-Time Notetaking from Keypoints with Large Language Models

### v1

@@ 8.2. NoTeeline supports a Wide Range of Notetaking Pattern and Associated Transition
Transitions. NoTeeline supports five types of transitions during note expansions. 1) Abbreviation Expansion: Converts abbreviations into their full form; 2) Detailed Information Extraction: extract missing information from transcript required for note expansion. 3) Personal Writing Style Imitation: mimic the writing pattern users usually use in their writeup; 4) Symbolic Interpretation: accurately decodes the symbols in user-specific writing style and integrates them into the expanded notes; 5) Typographical Error fix: corrects spelling mistakes or missing letters in the text.
8.3. Comparison of NoTeeline and Baseline
8.3.1. Workload.
Using the Shapiro-Wilk test of normality, the differences between baseline and experimental condition responses for three of the four NASA-TLX questions (Hart and Staveland, 1988) on workload significantly deviate from a Normal distribution. Hence, we compare the differences in workload through a Wilcoxon One-tailed Signed-Rank Test (Neuhäuser, 2011) with a Bonferroni-corrected (Weisstein, 2004) significance level of 0.0125 per condition. Participants report significantly lower results for Mental Demand, Effort, and Frustration ($p=0.0123,p=0.0024,p=0.0047$). We did not find statistically significance for higher Performance (p = 0.0803), this may be attributed to users’ general familiarity with notetaking, self-assessments of performance were already high in the baseline condition ($\mu=5.58$ in the baseline versus a slightly higher average of $\mu=6.25$ in the experimental condition on the 7-point Likert scale) and results may be subject to the ceiling effect (Hancock, 1989) (5 out of 12 participants gave the maximum 7 rating).
8.3.2. Usability.
Usability was measured via an adapted version of SUS (Brooke et al., 1996)) and yielded an overall average of 80.972, which corresponds to an A grade according to the Sauro–Lewis curved grading scale (Sauro and Lewis, 2016). We note that the average baseline SUS values are also high at 80.833. Thus, our experimental features do not seem to detract from the inherent usability of notetaking methods more familiar to users. Further, our baseline condition already provides salient advancements to basic notetaking by combining the video source and notetaking area with other AI-assisted features—all of which received generally positive comments from participants—into one area. This is salient as P2, P3, P4, P10, P12 brought up issues of screen space economy and the importance of a proper split screen setup during notetaking (e.g. P12 mentions advantages of our setup over typical split screen set-ups for notetaking, which pose difficulties for tab navigation).
8.3.3. Task Completion Time.
The differences in task completion times per participant across conditions pass the Shapiro-Wilk test of normality ($W=0.8665,p=0.0590$). Hence, we compare the differences in task completion times (seconds) through a parametric paired t-test. Participants report a significantly lower task completion time when using NoTeeline ($\mu=207.7,\sigma=17.25$) compared to the baseline ($\mu=322.36,\sigma=103.27$) ($t(11)=3.6,p=0.004$) (Figure 7(c)).
@@ 8.3. Comparison of NoTeeline and Baseline
In figure 8, we show participants rating for individual features in NoTeeline ($\mu=5.8,\sigma=1.15$) (7-point Likert scale, higher is better). Overall, NoTeeline was received positively by the users. Please see §8.7 for more detailed feedback on NoTeeline’s features.
8.4. Note quality of NoTeeline
Linguistic Property. We report the linguistic quality (in terms of grammaticality, non-redundancy, focus) of notes using GRUEN (Zhu and Bhat, 2020). The result (Table 5 in Appendix) indicate that the points written using NoTeeline ($\mu=0.84,\sigma=0.07$) are of higher quality (i.e: less grammatical less, less redundant and more focused) than the baseline ( $\mu=0.56,\sigma=0.17$) ($p<0.0001,t(11)=10.87$).
Correctness. To ensure our automatically expanded notes are indeed correctly extracting the information from the transcript and not hallucinating any misinformation (C2), we compute the factual correctness of the generated note using HHEM (Vectara, 2024; Zhang et al., 2023; Azamfirei et al., 2023). We noticed a Hallucination Rate = $6.8\%$ and a Factual Consistency Rate = $93.2\%$. This indicates that our generated notes are correct and consistent $93.2\%$ of the time.
User evaluation. We also collect the self-reported scores by the users. On average, users rated the expanded note, $\mu=6$ ($\sigma=1.044$) on a scale of 1 to 7 (higher is better) (Figure 8).
8.5. Consistency of Micronote and its Corresponding Expanded Note
To ensure our expanded notes are consistent with the main content of the micronotes, we measure the widely used sentence representation similarity, SBERT (Reimers and Gurevych, 2019; Ethayarajh, 2019; Gao et al., 2019). We find that they have a 58% similarity (avg. score = 0.58). We also examine the similarity of expanded note and the transcript and found it to be 42% (avg. score = 0.42). This results show that the expanded notes are more similar to micronotes than the transcript, indicating that our expanded notes prioritize the micronotes’ contents and do not merely replicate the video transcripts, which will be undesirable.
8.6. Consistency of Summary and Note
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12 avg.
Consistency of Micronote, transcript & corresponding expanded note (§8.5)
Micronote &
Expanded Note
0.6 0.41 0.73 0.75 0.63 0.58 0.55 0.6 0.48 0.43 0.57 0.63 0.58
Transcript &
Expanded Note
0.47 0.41 0.42 0.42 0.42 0.39 0.42 0.42 0.44 0.46 0.39 0.41 0.42
Consistency of Summary & Note in two conditions (§8.6)
NoTeeline 0.87 0.51 0.9 0.63 0.91 0.77 0.67 0.8 0.68 0.7 0.76 0.82 0.75
8.7. How NoTeeline Features Assist in Notetaking
8.7.1. Impact of Note Expansion.
@@ 8.7. How NoTeeline Features Assist in Notetaking
All 12 participants reviewed some aspect of the expanded note favorably. All critical feedback, present in four of the responses, focused on the potential for verbosity and redundancy in the expanded notes (refer to §8.4 for demonstration of how NoTeeline still improves upon the baseline for measures of redundancy). Still, nine reviews expressed appreciation for the quality/level of detail in the expanded note, and five reviews brought up that the expanded note captured details/context they would have otherwise missed, due to such reasons as cognitive and time constraints (e.g. P9: focusing on the key words is less mentally draining and would be helpful in keeping my attention span for a longer period of time’; P6: ‘there’s like a lot of things that I heard I probably wouldn’t have to have time to type it all out, but it was able to expand on everything’). Eight participants also specifically remarked that the expanded note reflected their writing style. P11 says ‘If my note was seven words or eight words then it went like one and a half lines if mine was small and then it was also like not trying to bring in too much content so that I felt it was trying to match with my style of what I’m trying to put’.
8.7.2. Impact of Generated Themes.
@@
11 out of 12 participants reviewed some aspect of the generated themes positively. P8 believes the impact of Order by Theme would become more evident in longer videos. P10 discusses how it would have greater impact in a video with more discursive content. In that regard, P11 expresses excitement about using the Order by Theme in such a context ‘where the conversation may not usually flow through on one topic’. Five reviews specifically point out the utility of the Order by Theme feature for aiding review, retrieval, and organization of content.
8.7.3. Impact of Cue Questions.
@@
11 out of 12 participants reviewed some aspect of the cue questions positively, with critical feedback from five participants tending to focus on increased cue question complexity. P9 identified that multiple choice question may not be applicable for ‘concepts [that] are more open to interpretation’. Six participants specifically brought up positive impressions of the cue questions aiding in the video/note review process, four participants mentioned that they viewed the cue questions as a way to identify potential shortcomings in their notes, and three participants highlighted perceived relevance of cue questions to the video content.
8.7.4. Impact of Summary.
@@
11 out of 12 participants reviewed some aspect of the summary feature positively. Four responses brought up satisfaction with the summary’s ability to capture the gist/important contexts from their notes. Three responses mentioned a preference for notes versus the generated summary for reviewing purposes: P1 prefers reviewing ‘small details’ while P3 and P6 find the length too long. However, opinion on length is not uniform, as five responses (P5, P7, P8, P10, P12) reflected positively on the length of the summary. Four responses described a desire for more formatting features, such as bolded or highlighted text.
8.7.5. General Reception of AI Assistance
@@
We conclude thematic analysis with a general review of reception to NoTeeline’s AI-assistance during notetaking. Across the user study, ten participants specifically addressed the relationship between their labor and the AI features in NoTeeline impacting their workflows. P1 speaks to the ‘division’ of labor in which they capture the ‘smaller details’ while the ‘machine… take[s] over the higher level summary.’ P1, P3-P4, P6-P8, P10-P12 speak to how the automated features distributed throughout NoTeeline provide enhancement or safeguard beyond their individual labor–whether capturing missed details/context, correcting typographic and grammatical errors or otherwise refining their writing, or surfacing unrecognized themes or shortcomings in their understanding of the material.
8.8. NoTeeline Enables Efficient Note-taking across Diverse Domains
During the freeform test (§7), Participants reported and used NoTeeline in a wide range of video categories. All 12 participants wished to continue using NoTeeline in future. P7 continued using NoTeeline for annotating interviews collected as part of their research even after our study session. P9 mentions the desire for integration with a current Learning Management System, Canvas, used at their university. Most participants thought NoTeeline would be most useful for long-form videos where there is a lot of information to capture rather than short videos which may not that essential to take notes on. The top categories they reported are: 1. Educational Video Content (×$\times 8$) 2. Professional Podcast and meetings (×$\times 3$) 3. Thematic analysis of research videos (×$\times 3$) 4. Personal Learning and Self-Education from how-to tutorials (×$\times 2$) 5. Financial and Investment Learning (×$\times 1$).
8.9. Error Analysis
Error in LLM generation. GPT-4 sometimes fails to generate the requested note. For example, P7 wrote a micronote for noting down her thoughts that is not directly discussed in the video. The model replied with - ‘Please provide the transcript related to the keypoint so I can assist you in creating a note.’. The occurrence of such error is 0.72% (4 out of 559 total calls).
@@ 9. Discussion
9.1. Implication of NoTeeline in Modern real-time Notetaking
Building upon the established concept of micronotes, NoTeeline supports real-time notetaking from videos. By leveraging large language models (LLMs), NoTeeline reduces the manual effort required for writing full notes, allowing users to focus on capturing core ideas rather than transcribing content verbose.
Our study shows that users were able to produce an equivalent or greater amount of information in their notes using fewer words and less time. This suggests that micronotes, when combined with NoTeeline’s automation, remain a highly effective method for real-time notetaking without requiring extensive user involvement or high precision. The system helps users concentrate on the material being discussed rather than the act of writing, addressing common pain points associated with the overwhelming complexity of existing note-taking tools.
While concerns may arise about potential disengagement with the material due to automation, we argue that NoTeeline enhances rather than detracts from the cognitive process. Users still initiate and guide the note-taking process, with NoTeeline simply assisting in transforming their initial input into structured notes. This elucidates the workflow as well as promotes deeper engagement by freeing users from the mechanical task of writing. Prior research on LLM-based writing assistants also report that automating the writing experience does not necessarily hinder the main objective but rather enables more refined thinking (Nakazawa et al., 2022; Göldi et al., 2024). Furthermore, NoTeeline incorporates a cue question feature inspired by the Cornell method, which has been shown in previous studies to improve long-term knowledge retention (Kornell, 2014; Evans and Shively, 2019; Yuniarti and Trisnawati, 2018).
9.2. Design Implication for Human-AI Collaborative Tools
In human-AI collaboration, trust plays a critical role in determining how effectively users engage with AI systems. In our Needfinding study, we likewise uncover user concerns about LLM hallucination and control in the domain of note-taking (e.g. P1’s and P2’s quotes in theme C2, §4.2). Research shows that trust in human-machine teaming is built through the reliability, transparency, and explainability of AI systems (Alhaji et al., 2020). By controlling the precise moments of note-taking instead of relegating the summarizing task entirely to note-taking software, we offer greater transparency and specificity in the note-taking process. This is further supported by the low hallucination rates we uncovered in Section 8.5. This suggests that ensuring users can trust the AI to expand micronotes in a way that aligns with their personal keypoints and style is essential for fostering productive collaboration. Additionally, the ability to modify or intervene in the AI’s output has also been shown to increase trust (Jeung and Huang, 2023). In our tool, we achieve this via our note editing feature and the involved nature of the general micronote expansion process.
Further, we uncover the benefits of designing for implicit guidance by aligning the processes of task completion and training processes. In our system, note-taking serves a dual purpose — simultaneously achieving the note-taking task and providing continuous guidance to the LLM to improve user experience in a positive feedback loop, reducing needing additional external signals.
@@ 10. Limitations and Future Work
Through NoTeeline’s testing phase, we uncovered potential areas for future research. First, we could further organize notes in multi-level, hierarchical list formats to improve the note review experience. Second, we can incorporate support for rich text formatting (i.e: text highlighting, support for different font styles and sizes), which can be implemented trivially by incorporating rich text formatting package (quilljs.com, 2019). Third, we can potentially reduce the response time of GPT-4 by parallel computing or pre-computing future operations. Some recent works streamline the model output to reduce the wait time (Lin et al., 2024), but we did not prefer it in NoTeeline. Our concern was that any inaccuracies in the streamlined output might complicate error correction, as the information is already presented to users, potentially compromising the integrity of the user experience. We can also consider the inclusion of non-textual note elements, such as images and drawings. Finally, as surfaced by participants multiple times throughout the user study, a promising direction can be evaluating the potential learning outcomes of NoTeeline, i.e: to what extent it can enhance academic performance and learning retention. Since we did not specifically consider academic contexts in NoTeeline, our current focus has been on its efficacy for swift, personalized notetaking (Lin et al., 2004). More generally, future work may further explore the complexities of AI-assisted notetaking in contrast to other writing domains (Bedington et al., 2024).
@@ 11. Conclusion
Active notetaking from videos remains challenging. While existing tools can automatically generate notes, they can not capture the stylistic nuances users care about. Our needfinding study reveals that users often do not prefer an automatically generated summary due to concerns about accuracy and lack of customization. On the other hand, NoTeeline leverages rich contextual information to generate more accurate and more personalized notes. Our user study reveals that NoTeeline helps write notes faster with $93.2\%$ correctness and achieves closer stylistic characteristics to users’ handwritten notes. We demonstrate the utility of NoTeeline in a variety of usage scenarios users reported and provide design implications for future works in human-AI collaboration.
@@
By helping users take notes faster, our tool has potential to make note-taking more accessible for individuals who might otherwise struggle due to reasons such as language fluency (e.g. P7’s comments about difficulties of being a non-native English speaker). While there exist general arguments about how over-reliance on AI-generated content may lead to passive engagement or diminish personal organizational skills, we show how our tool exists to amplify not override user engagement by generating content driven by direct user input. We expect this should also decrease the risk of AI biases that might occur in more fully generated content (e.g. information prioritized during note generation). Finally, while the mechanisms of our tool are not necessarily in themselves geared towards learning outcomes, facilitating faster and more extensive information capture–in a personalized way via a compact UI–has the possibility to help users organize, synthesize, and engage with content in ad hoc ways towards educational, professional, and personal use cases.

### v2

@@ 8.3.1. NoTeeline Supports Diverse Notetaking Patterns
Operations. We observed that NoTeeline supports five types of operations during note expansions. 1) Abbreviation Expansion: Converts abbreviations into their full form; 2) Detailed Information Extraction: extract missing information from transcript required for note expansion. 3) Personal Writing Style Imitation: mimic the writing pattern users usually use in their writeup; 4) Symbolic Interpretation: accurately decodes the symbols in user-specific writing style and integrates them into the expanded notes; 5) Typographical Error fix: corrects spelling mistakes or missing letters in the text.
8.3.2. NoTeeline Supports a Variety of Content Domains
During the freeform test (§7), Participants reported and used NoTeeline in a wide range of video categories. All 12 participants wished to continue using NoTeeline in the future. P7 reported that they continued using NoTeeline for annotating research interviews even after our study session. P9 mentions the desire for integration with a current Learning Management System, Canvas, used at their university. Most participants thought NoTeeline would be most useful for long-form videos where there is a lot of information to capture. The top categories they reported are: 1. Educational Video Content ($n=8$); 2. Professional Podcast and meetings ($n=3$); 3. Thematic analysis of research videos ($n=3$); 4. Personal Learning and Self-Education from how-to tutorials ($n=2$); 5. Financial and Investment Learning ($n=1$).
8.3.3. Comments on NoTeeline’s Individual Features
@@
In figure 10, we show participants rating for individual features in NoTeeline ($\mu=5.8,\sigma=1.15$) (7-point Likert scale, higher is better). Overall, NoTeeline was received positively by the users.
@@
Impact of Note Expansion. All 12 participants reviewed some aspect of the expanded note favorably. All critical feedback, present in four of the responses, focused on the potential for verbosity and redundancy in the expanded notes (refer to §8.2 for demonstration of how NoTeeline still improves upon the baseline for measures of redundancy). Still, nine reviews expressed appreciation for the quality/level of detail in the expanded note, and five reviews brought up that the expanded note captured details/context they would have otherwise missed, due to such reasons as cognitive and time constraints (e.g. P9: focusing on the keywords is less mentally draining and would be helpful in keeping my attention span for a longer period of time’; P6: ‘there’s like a lot of things that I heard I probably wouldn’t have to have time to type it all out, but it was able to expand on everything’). Eight participants also specifically remarked that the expanded note reflected their writing style. P11 says ‘If my note was seven words or eight words then it went like one and a half lines if mine was small and then it was also like not trying to bring in too much content so that I felt it was trying to match with my style of what I’m trying to put’.
@@
Impact of Generated Themes. 11 out of 12 participants reviewed some aspects of the generated themes positively. P8 believes the impact of Order by Theme would become more evident in longer videos. P10 discusses how it would have a greater impact in a video with more discursive content. In that regard, P11 expresses excitement about using the Order by Theme in such a context ‘where the conversation may not usually flow through on one topic’. Five reviews specifically point out the utility of the Order by Theme feature for aiding the review, retrieval, and organization of content.
@@
Impact of Cue Questions. 11 out of 12 participants reviewed some aspects of the cue questions positively, with critical feedback from five participants tending to focus on increased cue question complexity. P9 identified that multiple choice questions may not be applicable for ‘concepts [that] are more open to interpretation’. Six participants specifically brought up positive impressions of the cue questions aiding in the video/note review process, four participants mentioned that they viewed the cue questions as a way to identify potential shortcomings in their notes, and three participants highlighted the perceived relevance of cue questions to the video content.
@@
Impact of Summary. 11 out of 12 participants reviewed some aspect of the summary feature positively. Four responses brought up satisfaction with the summary’s ability to capture the gist/important contexts from their notes. Three responses mentioned a preference for notes versus the generated summary for reviewing purposes: P1 prefers reviewing ‘small details’ while P3 and P6 find the length too long. However, opinion on length is not uniform, as five responses (P5, P7, P8, P10, P12) reflected positively on the length of the summary. Four responses described a desire for more formatting features, such as bolded or highlighted text.
@@
General Reception of AI Assistance. We conclude the thematic analysis with a general review of reception to NoTeeline’s AI assistance during notetaking. Across the user study, ten participants specifically addressed the relationship between their labor and the AI features in NoTeeline impacting their workflows. P1 speaks to the ‘division’ of labor in which they capture the ‘smaller details’ while the ‘machine… take[s] over the higher level summary.’ P1, P3-P4, P6-P8, P10-P12 speak to how the automated features distributed throughout NoTeeline provide enhancement or safeguard beyond their individual labor–whether capturing missed details/context, correcting typographic and grammatical errors or otherwise refining their writing, or surfacing unrecognized themes or shortcomings in their understanding of the material.
@@ 9. Discussion
9.1. Summary of Findings
Building upon the established concept of micronotes, NoTeeline supports real-time notetaking from videos in a fast-paced environment. We found that NoTeeline reduces the manual effort required for writing full notes, allowing users to focus on capturing core ideas rather than the tedious efforts of transcribing content verbose. Our study shows that users were able to capture a large amount of content in their notes with significantly fewer words and less time compared to manual notes. Our study also reveals that users are able to use different note taking patterns while NoTeeline correctly decodes their micronote. All participants in the study preferred NoTeeline over the baseline and envisioned themselves using it in the future. By providing users the ability to write their own notes while enabling LLM assisted automated in a personalized way, NoTeeline achieves user trust and gives them a sense of agency.
9.2. Supporting Uninterrupted Active Notetaking Experience
While one might argue that spending more time to write individual notes is a better way to get a deeper engagement with the materials, and NoTeeline might seem to deviate users from writing effective notes, we argue that NoTeeline does not detract, but rather enhances the engagement in more meaningful ways.
First, users are still responsible for initiating which parts of the material to take notes for and what information they want to capture, with NoTeeline simply helping to transform their input into fuller structured notes more suitable for later review. In contrast to existing fully automated tools, users are actively involved in notetaking.
Second, NoTeeline reduces the disruption in watching videos caused when taking full notes (e.g., constantly having to pause the video) and freeing them from the mechanical task of writing. Thus, It helps maintain the uninterrupted flow of information.
Third, NoTeeline also allows deeper involvement in note analysis and revision stages. After the users have finished taking the notes, they can always come back to it and edit, reorganize, and revise based on their needs. Our system provides features to review notes by the cue question, to manually edit the automatically expanded notes, to jump to the location where the note was taken, and to organize the notes in their desired structure and themes. These features were inspired by the Cornell notetaking system with their long history of improving long-term knowledge retention (Kornell, 2014; Evans and Shively, 2019; Yuniarti and Trisnawati, 2018). All these factors altogether enable users to deeply engage with content in NoTeeline.
9.3. Insights for Human-AI Collaborative Notetaking Tools
Based on our findings from the design and evaluation of NoTeeline, we highlight three insights to inform the future design of human-AI collaborative notetaking systems. First, our findings reveal that the greatest benefit of NoTeeline is the affordance of the micro note and its expansion. Being able to continue writing notes within their usual workflow while enjoying LLM assistance sparks interest. The ability to toggle back and forth between the original and modified notes enhances their overall experience by providing control over the content, a critical factor in fostering user trust (Jeung and Huang, 2023). Future tools should support iterative processing, including automatic cleaning of grammatical errors, and offer both LLM-assisted and manual options for editing and organizing notes.
Second, collaborative notetaking systems should provide granular user control. In our needfinding study, users mention concerns about LLM hallucination and lack of control in AI-assisted tools (§4.2). Prior Research shows that trust in human-machine teaming can be positively impacted through transparency (Alhaji et al., 2020). By giving users the ability to control precise moments to take notes and writing styles to maintain – instead of relegating the generation task entirely to a software – NoTeeline can increase transparency in the notetaking process.
Finally, an AI-assisted system should leverage implicit guidance (Jannach et al., 2018) from users to regulate its generation quality. In NoTeeline, users would take micronotes as usual while implicitly guiding the system about what to focus on in the generated note. Such support for implicit guidance does not require additional effort from the user such as writing custom prompts for their individual requests every time.
@@ 10. Limitations and Future Work
While NoTeeline was generally positively received, there are some avenues for future work that we would like to explore. First, NoTeeline enables users to take notes as they progress through the content, following a linear notetaking workflow (Meng et al., 2016). It does not directly provide support for writing non-linear notetaking (though does provide support for non-linear organization of the notes by theme). Second, an interesting future extension could be to incorporate NoTeeline with a more visual notetaking tool such as VideoSticker (Cao et al., 2022), so that users can capture images and figures from videos within their notes.
While our study demonstrates that NoTeeline significantly reduces cognitive load and improves notetaking efficiency by minimizing writing effort and time, we did not measure the long-term effect of NoTeeline on information retention. The primary goal of this work was to evaluate the system’s ability to facilitate efficient note-taking, not to assess how the generated notes impact user learning outcomes. This follows standard practice in prior research, as seen in works like Rambler(Lin et al., 2024) and VideoSticker(Cao et al., 2022). Nonetheless, for future work, we could investigate how NoTeeline can influence information retention through a longitudinal study.
@@ 11. Conclusion
Active notetaking from videos remains challenging. While existing tools can automatically generate notes, they do not necessarily capture the details and stylistic nuances users care about. Our needfinding study reveals that users often do not prefer automatically generated notes due to concerns about accuracy and lack of customization. On the other hand, NoTeeline leverages rich contextual information to generate more accurate and more personalized notes. Our user study reveals that NoTeeline helps write notes faster with $93.2\%$ correctness and achieves closer stylistic characteristics to users’ handwritten notes. We demonstrate the utility of NoTeeline in a variety of usage scenarios users reported and provide design implications for future works in AI-assisted notetaking systems.
@@
By helping users take notes faster, NoTeeline has the potential to make notetaking more accessible for individuals who might otherwise struggle with factors such as language fluency (e.g., as demonstrated by P7’s comments about difficulties of being a non-native English speaker) and physical disabilities. While there exist arguments about whether over-reliance on AI-generated content may lead to passive engagement or diminish personal organizational skills, we show how our tool exists to amplify, not override, user engagement in the generation process. We also expect this to decrease the risk of AI biases that might otherwise occur in fully autogenerated notes without any sort of user engagement.